OpenStudy (anonymous):

A large cylinder contains a saline solution for science experiments. The cylinder has a height of 40 mm and a radius of 12 mm. Braddock pours the saline solution into test tubes that each hold 3014 mm3. How many test tubes can Braddock prepare? Use 3.14 to approximate pi, and express your answer as a whole number. _______

OpenStudy (anonymous):

@EclipsedStar

OpenStudy (anonymous):

@TheDream

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

@TheDream

OpenStudy (anonymous):

Basically the same thing as last time

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[V= \pi*r^2h\]

OpenStudy (anonymous):

V =pi * 12^2 times 40

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

kk i'm gonna work it out and i'll tell you when i'm done

OpenStudy (anonymous):

904.32 \[:)\]

OpenStudy (anonymous):

@TheDream

OpenStudy (anonymous):

12^2=144, 144*40 = 5760, 5760*3.14 = 18086.4

OpenStudy (anonymous):

That is the value for volume. To find the answer you then divide 18086.4/3014 = 6

OpenStudy (anonymous):

@TheDream

OpenStudy (anonymous):

ya sorry thanx again

OpenStudy (anonymous):

np